OpenStudy (anonymous):

Input in standard form the equation of the given line. The line through (0, -3) and (3, 0)

random231 (random231):

(y+3)=m(x) m=(-3-0)/(0-3)

OpenStudy (anonymous):

that doesn't help me :(

OpenStudy (texaschic101):

first find the slope using the slope formula : (y2 - y1) / (x2 - x1) (0,-3)..x1 = 0 and y1 = -3 (3,0)...x2 = 3 and y2 = 0 now we sub (0 - (-3) / (3 - 0) (0 + 3) / 3 3/3 = 1 so our slope is 1 now we use y = mx + b slope(m) = 1 you can use either of your points...(3,0)..x = 3 and y = 0 now we sub 0 = 1(3) + b 0 = 3 + b -3 = b our equation is : y = x - 3 but we need it in standard form Ax + By = C y = x - 3 -- subtract x from both sides -x + y = -3 -- multiply by -1 to make x positive x - y = 3 <== standard form

OpenStudy (triciaal):

slope of the line = change in y/corresponding change in x points (0, -3) (3,0) 0--3/3-0 = 3/3 = 1 using one of the points (3,0) and y = mx + b 0 = (1)(3) + b gives b = -3 y = x -3 standard form Ax + By = C rewrite to this form x - y = 3

OpenStudy (texaschic101):

great minds think alike :)